An Alka-Seltzer tablet consists of NaHCOs (sodium bicarbonate), citric acid (H3C&HsO; citric acid is triprotic), aspirin andHCI and NaHCOs starts, it will produce CO2 gas which will displaces water from the wash bottle into a graduated cylinder. AftT is the temperature of the gas in degrees Kelvin . Remembering that that the number of moles, n, is equal to the mass (g) di. In the procedure, you are asked to load as much sodium bicarbonate into the gelatin capsule as possible. Why is it advantagC. Calculate the density of CO(g) in g/L D. Calculate the molar mass of CO2 assuming a temperature ot 293.6 3. Using a sampleI Need the pre-lab Questions, please

An Alka-Seltzer tablet consists of NaHCOs (sodium bicarbonate), citric acid (H3C&HsO; citric acid is triprotic), aspirin and other inert ingredients. When dissolved in water, citric acid and NaHCO3 components react as follows: In the first part of the experiment, we will determine the mass of NaHCO in a tablet of Alka Seltzer. To do this, we will use a known mass of a powdered sample of Alka-Seltzer and determine the mass of CO2 gas which it generates. For this analysis to work, NaHCO must be the limiting reagent. Since this is not true for the components in the tablet, we will add HCl(aq). HCl(aq) is a strong acid while citric acid is a weak acid. With the two acids in the same solution, the NaHCO3 will first react with HCI and NaHCO will react with citric acid only if the HCI has been completely consumed. The amount of HCI we will use in this lab is in excess so that the citric acid will not react. The reaction between NaHCO3 and HCl is NaHCOs(aq) HCl(aq NaClaqH20) CO (g (eq 1) In order to determine the density of CO2 as well as the ideal gas constant (R), the lab procedure utilizes a gelatin capsule, test tube, wash bottle and 4 M HCl(aq). The gelatin capsule will hold a powdered Alka-Seltzer sample and this will be placed in a test tube with 4 M HCL The gelatin capsule takes 4-5 minutes to dissolve in 4 M HCl(aq) and this is sufficient time to perform weighing operations, place the test tube in the water filled wash bottle and close the lid, all of these prior to any reaction. When the gelatin capsule eventually dissolves the reaction between
HCI and NaHCOs starts, it will produce CO2 gas which will displaces water from the wash bottle into a graduated cylinder. After the reaction is completed, the outside of the test tube will be dried and the test tube weighed. The difference in mass is equal to the mass of carbon dioxide produced and the volume of the CO (g) is equal to the mL of water displaced from the wash bottle. In addition to these measurements, the temperature of CO2 inside the wash bottle and the room atmospheric pressure will be measured. The computation of mass of NaHCO in the Alka-Seltzer tablet follows the generalized scheme g CO2-mols CO2 mols NaHCO, g NaHco, % m ass NaHco, mass of NaHCO, in a tablet (tablet mass is -3.50 g) For example, if 0.102 g CO, is produced by 0.350 g of Alka-Seltzer sample, the calculations are: Lnr CO m lao u0u NacO9gNo 84.04 gNa 0.195 g NaHCO, 1 mol NaHCO, This 0.195 g represents the amount of NaHCOs in the initial 0.350 g Alka-Seltzer sample. This indicates that one tablet contains 100 (0.195 g / 0.350 g) -55.7 % NaHCO, Using this percentage, the mass of NaHCO in the original tablet is equal to Mass NaHCO, in tablet-0.557 x 3.50 g-1.95 g NaHCO, where the total weight of an Alka-Seltzer tablet is 3.50 g During the second part of the experiment we will determine of density of CO (g) and the determination of the molar mass of CO, using the Ideal Gas Law The ideal gas law is PV-nRT where . · . P is the pressure of the gas expressed in atmospheres V is the volume ofthe gas expressed in liters N is the number of moles of the gas
T is the temperature of the gas in degrees Kelvin . Remembering that that the number of moles, n, is equal to the mass (g) divided by the molar mass, M (g/mol). Therefore PV RT Isolating m and V onto one side of the equation produces m PMM V RT where m/V is the density of the gas (g/L) dRT Rearranging, the molar mass equation is According to this equation, in addition to the density, measurements of the temperature and pressure of the gas are required. The parameters on the right hand side of equation 11 will be measured as follows P: inside wash bottle)-P(CO)-P(atmosphere). We will measure the atmospheric pressure in the lab We will measure the temperature after the reaction has gone to completion by opening the lid slightly and inserting a temperature measurement device. We will calculate density from the measurements of the mass of CO2 and the volume of the water displaced from the wash bottle. T: d: Lastly, we will determine the ideal law constant (R). In the ideal gas law PV = nRT, the role of R is the proportionality constant. All proportionality constants must be determined experimentally. If P is measured in atmospheres, V is measured in L, T is measured in units of K and n is moles, the proportionality constant is known to have the value of R-0.08206 This value of R will be confirmed using a rearrangement of the ideal gas law where Latm Κ.mol PV
. In the procedure, you are asked to load as much sodium bicarbonate into the gelatin capsule as possible. Why is it advantageous to load a large sample mass compared to a small mass such as 0.050 g? In which case would you expect the error of your measurements of the amount of CO2 gas produced to be less? 2. Based on the procedure described in the lab, the data collected was as follows: Mass (before reaction): test tube HCIa)+stir bar +capsule 26.042g Mass (after reaction): test tube +HCla)+stir bar+capsule 25.783 g Volume of water displaced from the squirt bottle Temperature of the CO:(g CO2(g) pressure 144 mL 293.6 K 0.986 atm Calculate cach of the items below to three significant figures A. Calculate the mass of CO (g) B. Calculate the moles of CO (g)
C. Calculate the density of CO(g) in g/L D. Calculate the molar mass of CO2 assuming a temperature ot 293.6 3. Using a sample of 0.385 g of Alka-Seltzer, the mass of CO is found to be 0.122g. Calculate to three significant figures. A. The mass of NaHCOs that produced the 0.122 g of CO using the eq 1 from the introduction section. B. The % mass of NaHCO, in the original sample of Alka-Seltzer. C. The mass of NaHCO, in the tablet assuming the mass of the tablet is 3.50 g.
Answer #1

1. As you load large mass of sample, more CO2 is produced and there will be less error in calculations.

2. (a)

Mass of CO2 = (26.042 g - 25.783 g ) = 0.259 g

(b) Moles of CO2:

n = P *V / R * T = 0.986 atm * 0.144 L / ( 0.08206 Lit atm mole-1 K-1 * 293.6 K ) = 0.00589 mole

(c) Density of CO2 = mass/ Volume = 0.259 g / (0.144 L) = 1.80 g/ L

(d) Molar mass of CO2 (M) = d*R*T / P = (1.80 g / lit) * 0.08206 Lit atm mole-1 K-1* 293.6 K/ 0.986 atm = 44.0 g/ mole

3. (A)

Moles of CO2 produced = mass/ molar mass of CO2 = 0.122 g / (44 g / mole) = 0.00277 mole

Moles of NaHCO3 required to produce the CO2 = 0.00277 mole

Mass of HaHCO3 = moles * Molar mass of NaHCO3 = 0.00277 mole * (84 g/ mole) = 0.233 g

(B) % mass of NaHCO3 = 0.233 g * 100/ 0.385 g % = 60.5%

(C) Mass of NaHCO3 in 3.50 g of the tablet = 3.50 g * (60.5/100) = 2.12 g